OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through the given point and is perpendicular to the graph of the equation. 1.(-5,2) y=1/2x-3

OpenStudy (texaschic101):

y = 1/2x - 3 (the slope here is 1/2, but for a perpendicular line, you want the negative reciprocal of that slope. To find the negative reciprocal, you " flip " the slope and change the sign. The negative reciprocal of 1/2 is -2/1 or just -2. That is the slope we need. Now we will use this formula : y - y1 = m(x - x1) using slope(m) = -2 (-5,2) x1 = -5 and y1 = 2 Now lets sub.. y - 2 = -2(x - (-5) y - 2 = -2(x + 5) y - 2 = -2x - 10 y = -2x - 10 + 2 y = -2x - 8 ==>slope intercept form